# Cumulants in rectangular finite free probability and beta-deformed singular values

### v1

@@ 3.3 In terms of Łukasiewicz paths
$m_{2k}=\sum_{P\in\mathbf{L}^{\textrm{odd}}(2k)}\prod_{s\geq 0}{(u+s)^{\#\textrm{down steps of $P$ from height }(2s+1)}}\\ \cdot\prod_{s\geq 1}{(t+s)^{\#\textrm{down steps of $P$ from height }(2s)}}\prod_{s\geq 1}{\kappa_{2s}^{\#\textrm{up steps }(1,2s-1)\textrm{ of $P$}}}.$ (29)
In other words, for any odd Łukasiewicz path $P$ of length $2k$, associate to each of its steps a weight according to the following rules. To each up step $(1,2s-1)$, associate the weight $\kappa_{2s}$ and to each down step $(1,-1)$ from height $(j+1)$ to height $j$ the weight $(u+s)$, if $j=2s$ is even, and the weight $(t+s)$, if $j=2s-1$ is odd. [^6] By definition, there are no horizontal steps in an odd Łukasiewicz path. Finally, associate to $P$ the product of all weights of its steps and denote it by $w_{t,u}(P)$. Then $m_{2k}$ is the sum of weights $w_{t,u}(P)$, as $P$ ranges over all odd Łukasiewicz paths of length $2k$.
This result is equivalent to [Xu-23, Thm. 5.5]. Indeed, that result gives an equivalent formula in terms of noncrossing even set partitions, which relate to odd Łukasiewicz paths as explained by Cor. 2.3. The advantage of our theorem is that the formula in terms of Łukasiewicz paths is more natural and also bears resemblance to universal formulas in the related discrete context of high temperature beta-partitions [CDM-23, CD-24+].
@@
where the sum is over “words” $a_{2k}\cdots a_{2}a_{1}$ of length $2k$ and each $a_{i}$ is either the operator $\partial_{t,u}$ or $*_{g}$, but the last “letter” is $a_{1}=*_{g}$. Since $g(z)=\sum_{\ell\geq 1}{\kappa_{2\ell}z^{2\ell-1}}$, this sum can be further refined and written as:
@@
where each $b_{i}$ is either $\partial_{t,u}$ or $*_{\kappa_{2\ell}z^{2\ell-1}}$ (operator of multiplication by $\kappa_{2\ell}z^{2\ell-1}$), for some $\ell\in\mathbb{Z}_{\geq 1}$, and $b_{1}$ is of the latter type.
@@
$\bullet$ each $c_{i}$ is either $\partial_{t,u}$ or $*_{\kappa_{2\ell}z^{2\ell-1}}$, for some $\ell\in\mathbb{Z}_{\geq 1}$;
@@
with certain edge-weights according to the following procedure. If $c_{1}=*_{\kappa_{2\ell}z^{2\ell-1}}$, then set $w_{1}:=(1,2\ell-1)$ and put an edge-weight of $\kappa_{2\ell}$ on $(w_{0}\to w_{1})$. In general, if $w_{1},\dots,w_{j-1}$ have already been chosen and $c_{j}=*_{\kappa_{2\ell}z^{2\ell-1}}$, then set $w_{j}:=w_{j-1}+(1,2\ell-1)$ and assign an edge-weight of $\kappa_{2\ell}$ on $(w_{j-1}\to w_{j})$. On the other hand, if $c_{j}=\partial_{t,u}$, then set $w_{j}:=w_{j-1}+(1,-1)$. The weight of the edge $(w_{j-1}\to w_{j})$ in this case will depend on the parity of $\deg(c_{j-1})+\cdots+\deg(c_{1})$: if this is even and equal to $2s$, assign the weight $(t+s)$ to the edge, whereas if it is odd and equal to $12s+$ 1, assign the weight $(u+s)$ to the edge. By induction, one shows that for $j=1,2,\dots,2k$, the suffix $c_{j}\cdots c_{1}(1)$ equals $r_{j}z^{s_{j}}$, where $s_{j}=\deg(c_{1})+\cdots+\deg(c_{j})$ and $r_{j}$ is the multiplication of edge-weights of $(w_{0}\to w_{1}),\dots,(w_{j-1}\to w_{j})$. In particular, the term $c_{2k}\cdots c_{1}(1)$ in (31) equals the weight of the whole constructed Łukasiewicz path; this ends the proof. ∎
Figure 1: The four odd Łukasiewicz paths of length $6$ with two up steps $(1,3)$ and six down steps $(1,-1)$. All edge-weights and weights of the Łukasiewicz paths are also shown.
@@
Let us verify that our Thm. 3.6 correctly predicts the coefficient of $\kappa_{4}^{2}$ in $m_{8}$. Indeed, a term in the sum (29) is a multiple of $\kappa_{4}^{2}$ if it corresponds to some odd Łukasiewicz path with two up steps $(1,3)$ and six down steps $(1,-1)$. There are exactly four such Łukasiewicz paths; they are exhibited, together with their step-weights, in Fig. 1. The sum of the four weights of the Łukasiewicz paths displayed is indeed $u(t+1)(u+1)(t+u+3)(2tu+3t+3u+6)\kappa_{4}^{2}$, matching (32).
@@ 4.2 Symmetric additive convolution
$\mathbf{c}=\mathbf{a}*\mathbf{b}=\left(\underbrace{\frac{c_{1}}{(-d)_{1}},\frac{c_{2}}{(-d)_{2}},\dots,\frac{c_{d}}{(-d)_{d}}}_{\text{first }d\text{ entries of $a*b$}},\underbrace{*,*,\dots,*}_{\text{ entries }(d+1),\dots,2d},\underbrace{0,0,0,\dots}_{\text{ entries $>2d$ vanish}}\right).$
@@
$K_{\ell}[p]:=\ell\cdot[z^{\ell}]\,\ln\left(1+\sum_{i=1}^{d}{\frac{a_{i}}{(-d)_{i}}z^{i}}\right),\quad\ell=1,\dots,d.$ (41)
As a consequence of Lem. 4.1, we obtain the following corollary that states that finite free cumulants linearize the symmetric additive convolution.
@@
The quantities $-K_{\ell}[p]\cdot d^{\ell-1}$, $\ell=1,\dots,d$, coincide with the version of finite free cumulants considered in [AP-18], denoted there by $\kappa^{p}_{\ell}$. The previous corollary is therefore equivalent to Prop. 3.6 in that paper, though our proof is different.
@@ 4.3 Rectangular convolution
For any monic polynomial $p(x)$ of degree $d$ and any $n\in\mathbb{Z}_{\geq 0}$, define its $\boldsymbol{(n,d)}$-rectangular (finite free) cumulants $K^{n,d}_{2}[p],K^{n,d}_{4}[p],\dots,K^{n,d}_{2d}[p]$ as the quantities
@@
In the special case $n=0$, the $(n\!=\!0,\,d)$-rectangular convolution $\boxplus^{n=0}_{d}$ is called the asymmetric additive convolution in [MSS-22]. In this case, our $(n\!=\!0,\,d)$-rectangular cumulants from Def. 4.3 reduce to the asymmetric cumulants considered in [CS-24].
@@ 5.1 Limits of (n,d)-rectangular cumulants
The theorem and corollary in this subsection can be deduced from results on the rectangular finite $R$-transform, obtained in [Gri-24], but the proofs are different.
@@
“If direction”: Limits of cumulants implies limits of moments. Let $f_{d}(x)=x^{d}+\sum_{j=1}^{d}{a_{2j}^{(d)}x^{d-j}}$ be an arbitrary monic polynomial of degree $d$ with only nonnegative real roots. If the even moments of $\widetilde{\mu}[f_{d}]$ are denoted $\widetilde{m}_{2}[f_{d}],\widetilde{m}_{4}[f_{d}],\cdots$, then
@@
Note that equations (45)–(46) are precisely a specialization of (9)–(10) where, in particular, $t=-d$ and $u=-d-n$. Then, by Thm. 3.6, we have the following formula, for all $k\in\mathbb{Z}_{\geq 1}$:
$\widetilde{m}_{2k}[f_{d}]=\sum_{P\in\mathbf{L}^{\textrm{odd}}(2k)}\prod_{s\geq 0}{(-d-n+s)^{\#\textrm{down steps of $P$ from height }(2s+1)}}\\ \cdot\prod_{s\geq 1}{(-d+s)^{\#\textrm{down steps of $P$ from height }(2s)}}\prod_{s\geq 1}{\left[\frac{\widetilde{\kappa}_{2s}^{n,d}[f_{d}]}{(-d)^{2s-1}q^{s}}\right]^{\#\textrm{up steps }(1,2s-1)\textrm{ of $P$}}}.$ (47)
@@
follows from the fact that any Łukasiewicz path begins and ends at the same height $y=0$ (so the distance traveled going up must equal the distance traveled going down) and odd Łukasiewicz paths only have up steps of the form $(1,2s-1)$, for some $s\in\mathbb{Z}_{\geq 1}$.
@@
$\widetilde{m}_{2k}[f_{d}]=q^{-k}\cdot\sum_{P\in\mathbf{L}^{\textrm{odd}}(2k)}\prod_{s=0}^{k-1}{\left(1+\frac{n}{d}-\frac{s}{d}\right)^{\#\textrm{down stepsof $P$ from height }(2s+1)}}\\ \cdot\prod_{s=1}^{k-1}{\left(1-\frac{s}{d}\right)^{\#\textrm{down steps of $P$from height }(2s)}}\prod_{s=1}^{k}{\big{(}\widetilde{\kappa}^{n,d}_{2s}[f_{d}]\big{)}^{\#\textrm{up steps }(1,2s-1)\textrm{ of $P$}}},$ (50)
where we also changed infinite products to finite ones (e.g. $\prod_{s\geq 0}$ to $\prod_{s=0}^{k-1}$) by using the fact that any $P\in\mathbf{L}^{\textrm{odd}}(2k)$ has no vertices above height $(2k-1)$ or up steps $(1,2s-1)$ with $s>k$.
@@
$P_{2k}^{\,\widetilde{\kappa}\mapsto\widetilde{m}}(x_{2},x_{4},\dots,x_{2k}\,|\,d,n,q):=q^{-k}\cdot\sum_{P\in\mathbf{L}^{\textrm{odd}}(2k)}\prod_{s=0}^{k-1}{\left(1+\frac{n}{d}-\frac{s}{d}\right)^{\#\textrm{down steps of $P$ from height}\,2s+1}}\\ \cdot\prod_{s=1}^{k-1}{\left(1-\frac{s}{d}\right)^{\#\textrm{down steps of $P$from height}\,2s}}\prod_{s=1}^{k}{x_{2s}^{\#\textrm{up steps }(1,2s-1)\textrm{ of $P$}}},$ (51)
@@
$\lim_{d\to\infty,\,1+\frac{n}{d}\to q}{P_{2k}^{\,\widetilde{\kappa}\mapsto\widetilde{m}}(x_{2},\dots,x_{2k}\,|\,d,n,q)}\\ =q^{-k}\cdot\sum_{P\in\mathbf{L}^{\textrm{odd}}(2k)}{q^{\sum_{s=0}^{k-1}\#\textrm{down steps of $P$ from height }(2s+1)}}\prod_{s=1}^{k}{x_{2s}^{\ \#\textrm{up steps }(1,2s-1)\textrm{ of $P$}}}.$ (54)
@@
=#$=\#\,\textrm{even numbers in $[2k]$ not being the first in their blocks in $\pi$}$
$=k-\#\,\textrm{even numbers in $[2k]$ being the first in their blocks in $\pi$}$
@@
+some polynomial in the variables $+\textrm{some polynomial in the variables }x_{2},\dots,x_{2k-2},$
@@
[…]
@@ 5.2 Proof of the Main Result III
We learned about this in the case $n=0$ from [MSS-22], who attributed it to Mirabelli [Mir-21].
@@
$[z^{2\ell}]\ln\left(1+\sum_{k=1}^{d}{\frac{(-s^{2})^{k}}{n^{k}k!}z^{2k}}\right)=[z^{2\ell}]\ln\left(1+\sum_{k=1}^{\infty}{\frac{(-s^{2})^{k}}{n^{k}k!}z^{2k}}\right)\\ =[z^{2\ell}]\ln\left(\exp(-s^{2}z^{2}/n)\right)=[z^{2\ell}](-s^{2}z^{2}/n)=-\delta_{\ell,1}\cdot\frac{s^{2}}{n},$
@@
[…]
@@
[…]
@@
In this section, [^7] Since $q$ has already appeared as the parameter for the theory of rectangular free probability, we use the bold $\mathbf{q}$ in this section to denote the “quantum” deformation parameter. we consider a $\mathbf{q}$-analogue of the Main Result I from the introduction; in order to state it, we need some classical notations from $\mathbf{q}$-analysis, e.g. from [GR-11]. The $\mathbf{q}$-numbers are $[n]_{\mathbf{q}}=(1-\mathbf{q}^{n})/(1-\mathbf{q})$, $n\in\mathbb{Z}_{\geq 1}$, while $[0]_{\mathbf{q}}=1$; the $\mathbf{q}$-factorial is $n!_{\mathbf{q}}=[1]_{\mathbf{q}}[2]_{\mathbf{q}}\cdots[n]_{\mathbf{q}}$, for $n\in\mathbb{Z}_{\geq 1}$, while $10!_{\mathbf{q}}=$ 1; the $q$-exponential function is $\exp_{\mathbf{q}}(x)=\sum_{n=0}^{\infty}{x^{n}/n!_{\mathbf{q}}}$.
@@
In the relations above, we used the $\mathbf{q}$-numbers $[\ell]_{\mathbf{q}}=\frac{1-\mathbf{q}^{\ell}}{1-\mathbf{q}}$, $[k]_{\mathbf{q}}=\frac{1-\mathbf{q}^{k}}{1-\mathbf{q}}$, but $\mathbf{q}^{-\frac{1}{2}}$-composition (namely, definitions (78)–(79) should be employed with the parameter $\mathbf{q}^{-\frac{1}{2}}$ instead of $\mathbf{q}$) for the expressions of the form $\exp_{\mathbf{q}^{-\frac{1}{2}}}[f(z)]$ on the left hand sides of (81)–(82).
@@
    where $g(z):=\sum_{\ell=1}^{\infty}{\kappa_{2\ell}z^{2\ell-1}}$; $T_{\mathbf{q}^{-\frac{1}{2}}}$ is the shift operator $T_{\mathbf{q}^{-\frac{1}{2}}}f(z):=f\big{(}\mathbf{q}^{-\frac{1}{2}}z\big{)}$; $[z^{0}]f(z)$ is the constant term of the power series $f(z)\in\mathbb{R}[[z]]$; and finally, $\Delta_{\mathbf{q};t,u}$ is the linear operator on $\mathbb{R}[[z]]$ defined uniquely by its action on monomials:
@@
    $m_{2k}=\sum_{P\in\mathbf{L}^{\textrm{odd}}(2k)}\mathbf{q}^{\#\textrm{down steps of $P$}+\frac{1}{2}(\#\textrm{up steps of $P$}-\textrm{sum of heights ofvertices of $P$})}\\ \cdot\prod_{s\geq 0}{\Big{(}(1-u\mathbf{q}^{s})/(1-\mathbf{q})\Big{)}^{\#\textrm{down steps of $P$ from height }(2s+1)}}\\ \cdot\prod_{s\geq 1}{\Big{(}(1-t\mathbf{q}^{s})/(1-\mathbf{q})\Big{)}^{\#\textrm{down steps of $P$ from height }(2s)}}\prod_{s\geq 1}{\kappa_{2s}^{\#\textrm{up steps }(1,2s-1)\textrm{ of $P$}}}.$ (85)
@@
$\begin{gathered}\partial_{\mathbf{q};t,u}z^{2n+1}:=\frac{1-u\mathbf{q}^{-\frac{1}{2}}}{\mathbf{q}^{-1}-1}z^{2n},\quad n\in\mathbb{Z}_{\geq 0},\\ \partial_{\mathbf{q};t,u}z^{2n}:=\frac{1-t}{\mathbf{q}^{-1}-1}z^{2n-1},\quad n\in\mathbb{Z}_{\geq 1},\qquad\partial_{\mathbf{q};t,u}(1):=0.\end{gathered}$ (86)

### v3

@@ 3.3 In terms of Łukasiewicz paths
[…]
In other words, for any odd Łukasiewicz path $P$ of length $2k$, associate to each of its steps a weight according to the following rules. To each up step (1,2s1)(1,2s-1), associate the weight $\kappa_{2s}$ and to each down step (1,1)(1,-1) from height $(j+1)$ to height $j$ the weight $(u+s)$, if $j=2s$ is even, and the weight $(t+s)$, if j=2s1j=2s-1 is odd. [^6] By definition, there are no horizontal steps in an odd Łukasiewicz path. Finally, associate to $P$ the product of all weights of its steps and denote it by $w_{t,u}(P)$. Then $m_{2k}$ is the sum of weights $w_{t,u}(P)$, as $P$ ranges over all odd Łukasiewicz paths of length $2k$.
This result is equivalent to [Xu-23, Thm. 5.5]. Indeed, that result gives an equivalent formula in terms of noncrossing even set partitions, which relate to odd Łukasiewicz paths as explained by Cor. 2.3. The advantage of our theorem is that the formula in terms of Łukasiewicz paths is more natural and also bears resemblance to universal formulas in the related discrete context of high temperature beta-partitions [CDM-23, CD-25a, CD-25b].
@@
where the sum is over “words” $a_{2k}\cdots a_{2}a_{1}$ of length $2k$ and each $a_{i}$ is either the operator $\partial_{t,u}$ or $_{g}$, but the last “letter” is $a_{1}=*_{g}$. Since $g(z)=\sum_{\ell\geq 1}{\kappa_{2\ell}z^{2\ell-1}}$, this sum can be further refined and written as:
@@
where each $b_{i}$ is either $\partial_{t,u}$ or $_{\kappa_{2\ell}z^{2\ell-1}}$ (operator of multiplication by $\kappa_{2\ell}z^{2\ell-1}$), for some $\ell\in\mathbb{Z}_{\geq 1}$, and $b_{1}$ is of the latter type.
@@
$\bullet$ each $c_{i}$ is either $\partial_{t,u}$ or $_{\kappa_{2\ell}z^{2\ell-1}}$, for some $\ell\in\mathbb{Z}_{\geq 1}$;
@@
with certain edge-weights according to the following procedure. If $c_{1}=*_{\kappa_{2\ell}z^{2\ell-1}}$, then set $w_{1}:=(1,2\ell-1)$ and put an edge-weight of $\kappa_{2\ell}$ on $(w_{0}\to w_{1})$. In general, if $w_{1},\dots,w_{j-1}$ have already been chosen and $c_{j}=*_{\kappa_{2\ell}z^{2\ell-1}}$, then set $w_{j}:=w_{j-1}+(1,2\ell-1)$ and assign an edge-weight of $\kappa_{2\ell}$ on $(w_{j-1}\to w_{j})$. On the other hand, if $c_{j}=\partial_{t,u}$, then set $w_{j}:=w_{j-1}+(1,-1)$. The weight of the edge $(w_{j-1}\to w_{j})$ in this case will depend on the parity of $\deg(c_{j-1})+\cdots+\deg(c_{1})$: if this is even and equal to $2s$, assign the weight $(t+s)$ to the edge, whereas if it is odd and equal to $2s+1$, assign the weight $(u+s)$ to the edge. By induction, one shows that for $j=1,2,\dots,2k$, the suffix $c_{j}\cdots c_{1}(1)$ equals $r_{j}z^{s_{j}}$, where $s_{j}=\deg(c_{1})+\cdots+\deg(c_{j})$ and $r_{j}$ is the multiplication of edge-weights of $(w_{0}\to w_{1}),\dots,(w_{j-1}\to w_{j})$. In particular, the term $c_{2k}\cdots c_{1}(1)$ in (31) equals the weight of the whole constructed Łukasiewicz path; this ends the proof. ∎
Figure 1: The four odd Łukasiewicz paths of length 66 with two up steps $(1,3)$ and six down steps (1,1)(1,-1). All edge-weights and weights of the Łukasiewicz paths are also shown.
@@
Let us verify that our Thm. 3.6 correctly predicts the coefficient of $\kappa_{4}^{2}$ in $m_{8}$. Indeed, a term in the sum (29) is a multiple of $\kappa_{4}^{2}$ if it corresponds to some odd Łukasiewicz path with two up steps $(1,3)$ and six down steps (1,1)(1,-1). There are exactly four such Łukasiewicz paths; they are exhibited, together with their step-weights, in Fig. 1. The sum of the four weights of the Łukasiewicz paths displayed is indeed $u(t+1)(u+1)(t+u+3)(2tu+3t+3u+6)\kappa_{4}^{2}$, matching (32).
@@ 4.2 Symmetric additive convolution
[…]
@@
$K_{\ell}[p]:=\ell\cdot[z^{\ell}]\,\ln\left(1+\sum_{i=1}^{d}{\frac{a_{i}}{(-d)_{i}}z^{i}}\right),\quad\ell=1,\dots,d,$ (41)
where $z^{\ell}]f(z)$ denotes the coefficient of $z^{\ell}$ of any power series $f(z)\in\mathbb{R}[[z]]$. As a consequence of Lem. 4.1, we obtain the following corollary that states that finite free cumulants linearize the symmetric additive convolution.
@@
The quantities $K_{\ell}[p]\cdot d^{\ell-1}$, $\ell=1,\dots,d$, coincide with the version of finite free cumulants considered in [AP-18], denoted there by $\kappa^{p}_{\ell}$. The previous corollary is therefore equivalent to Prop. 3.6 in that paper, though our proof is different.
@@ 4.3 Rectangular convolution
For any monic polynomial $p(x)=x^{d}+\sum_{i=1}^{d}{x^{d-i}a_{2i}}$ of degree $d$ and any $n\in\mathbb{Z}_{\geq 0}$, define its $\boldsymbol{(n,d)}$-rectangular (finite free) cumulants $K^{n,d}_{2}[p],K^{n,d}_{4}[p],\dots,K^{n,d}_{2d}[p]$ as the quantities
@@
In the special case $n=0$, the (n=0,d)(n\!=\!0,\,d)-rectangular convolution $\boxplus^{n=0}_{d}$ is called the asymmetric additive convolution in [MSS-22]. In this case, our (n=0,d)(n\!=\!0,\,d)-rectangular cumulants from Def. 4.3 reduce to the asymmetric cumulants considered in [CS-24].
@@ 5.1 Limits of (n,d)-rectangular cumulants
The theorem and corollary in this subsection can be deduced from results on the rectangular finite $R$-transform, obtained in [Gri-24], but the proofs are different. We recall that for a polynomial $p(x)$ with nonnegative real roots, its symmetric empirical root distribution $\widetilde{\mu}[p]$ is defined in Eqn. (4).
@@
“If direction”: Limits of cumulants implies limits of moments. Let $f_{d}(x)=x^{d}+\sum_{j=1}^{d}{a_{2j}^{(d)}x^{d-j}}$ be an arbitrary monic polynomial of degree $d$ with only nonnegative real roots. If the even moments of $\widetilde{\mu}[f_{d}]$ are denoted $\widetilde{m}_{2}[f_{d}],\widetilde{m}_{4}[f_{d}],\cdots$, then we claim that
@@
In fact, if we denote the roots of $f_{d}(x)$ by $(\alpha_{1}^{(d)})^{2},\dots,(\alpha_{d}^{(d)})^{2}$, then $f_{d}(x)=x^{d}+\sum_{j=1}^{d}{a_{2j}^{(d)}x^{d-j}}=\prod_{k=1}^{d}{\big(x-(\alpha_{k}^{(d)})^{2}\big)}$. By multiplying both sides by $x^{-d}$ and doing the change of variables $x=z^{-2}$, we get
$1+\sum_{j=1}^{d}{a_{2j}^{(d)}z^{2j}}$ $=\prod_{i=1}^{d}{\Big(1-(\alpha_{i}^{(d)})^{2}z^{2}\Big)}=\prod_{i=1}^{d}{\big(1-\alpha_{i}^{(d)}z\big)\big(1+\alpha_{i}^{(d)}z\big)}$
$=\exp\left\{\sum_{i=1}^{d}\left(\ln\big(1-\alpha_{i}^{(d)}z\big)+\ln\big(1+\alpha_{i}^{(d)}z\big)\right)\right\}$
$=\exp\left\{-\sum_{i=1}^{d}\sum_{\ell=1}^{\infty}\frac{\Big(\big(\alpha_{i}^{(d)}\big)^{\ell}+\big(-\alpha_{i}^{(d)}\big)^{\ell}\Big)}{\ell}z^{\ell}\right\}$
$=\exp\left\{-d\sum_{k=1}^{\infty}\,\frac{\sum_{i=1}^{d}\frac{1}{2d}\Big(\big(\alpha_{i}^{(d)}\big)^{2k}+\big(-\alpha_{i}^{(d)}\big)^{2k}\Big)}{k}z^{2k}\right\},$
which immediately leads to the equality in Eqn. (45); the last equality in the chain of equations above follows because $\big(\alpha_{i}^{(d)}\big)^{\ell}+\big(-\alpha_{i}^{(d)}\big)^{\ell}=0$, if $\ell$ is odd, so the sum over $\ell\in\mathbb{Z}_{\geq 1}$ can be restricted to the positive even integers by the change of variables $\ell=2k$.
@@
Note that equations (45)–(46) are precisely a specialization of (9)–(10) where, in particular, t=dt=-d and u=dnu=-d-n. Then, by Thm. 3.6, we have the following formula, for all $k\in\mathbb{Z}_{\geq 1}$:
[…]
@@
follows from the fact that any Łukasiewicz path begins and ends at the same height $y=0$ (so the distance traveled going up must equal the distance traveled going down) and odd Łukasiewicz paths only have up steps of the form (1,2s1)(1,2s-1), for some $s\in\mathbb{Z}_{\geq 1}$.
@@
[…]
where we also changed infinite products to finite ones (e.g. $\prod_{s\geq 0}$ to $\prod_{s=0}^{k-1}$) by using the fact that any $P\in\mathbf{L}^{\textrm{odd}}(2k)$ has no vertices above height (2k1)(2k-1) or up steps (1,2s1)(1,2s-1) with $s>k$.
@@
[…]
@@
[…]
@@
[…]
@@
$+\textrm{some polynomial in the variables }x_{2},\dots,x_{2k-2},$
@@
$\lim_{d\to\infty,\,1+\frac{n}{d}\to q}{\widetilde{\kappa}^{n,d}_{2\ell}\big[p_{d}\boxplus^{n}_{d}r_{d}\big]}=\lim_{d\to\infty,\,1+\frac{n}{d}\to q}{\Big(\widetilde{\kappa}^{n,d}_{2\ell}[p_{d}]+\widetilde{\kappa}^{n,d}_{2\ell}[r_{d}]\Big)}\\ =\kappa_{2\ell}^{q}[\mu]+\kappa_{2\ell}^{q}[\nu]=\kappa_{2\ell}^{q}[\mu\boxplus_{q}\nu],$ (69)
@@ 5.2 Proof of the Main Result III
$z^{2\ell}]\ln\left(1+\sum_{k=1}^{d}{\frac{(-s^{2})^{k}}{n^{k}k!}z^{2k}}\right)=[z^{2\ell}]\ln\left(1+\sum_{k=1}^{\infty}{\frac{(-s^{2})^{k}}{n^{k}k!}z^{2k}}\right)\\ =[z^{2\ell}]\ln\left(\exp(-s^{2}z^{2}/n)\right)=[z^{2\ell}](-s^{2}z^{2}/n)=-\delta_{\ell,1}\cdot\frac{s^{2}}{n},$
@@
$\lim_{d\to\infty,\,1+\frac{n}{d}\to q}{(-d)^{2\ell-1}K^{n,d}_{2\ell}\left[\exp\left(-\frac{s^{2}}{n}x^{-n}Dx^{n+1}D\right)p_{d}\right]}\\ =\lim_{d\to\infty,\,1+\frac{n}{d}\to q}{(-d)^{2\ell-1}\left(K^{n,d}_{2\ell}[p_{d}]-\delta_{\ell,1}\cdot\frac{s^{2}}{n}\right)}=q^{-\ell}\kappa_{2\ell}^{q}[\mu]+\delta_{\ell,1}\cdot\frac{s^{2}}{q-1},$ (77)
@@
$\lim_{d\to\infty,\,1+\frac{n}{d}\to q}{(-d)^{2\ell-1}K^{n,d}_{2\ell}\left[\exp\left(-\frac{s^{2}}{n}x^{-n}Dx^{n+1}D\right)p_{d}\right]}$ $=q^{-\ell}\Big(\kappa_{2\ell}^{q}[\mu]+\kappa^{q}_{2\ell}\Big[\lambda_{qs^{2}/(q-1)}^{(q)}\Big]\Big)$
@@
Remark 2.
The theory of the rectangular finite free convolution is related to the classical Laguerre orthogonal polynomials. Indeed, recall that the Laguerre polynomial $L_{d}^{(\alpha)}(x)$ with parameter $\alpha>-1$ and degree $d$ is
$L_{d}^{(\alpha)}(x)=\sum_{k=0}^{d}{{d+\alpha\choose d-k}\frac{(-x)^{k}}{k!}}.$
Then the polynomial $r_{n,d}(x)$ in Eqn. (75) inside the proof above can be expressed in terms of the Laguerre polynomial $L_{d}^{(n)}(x)$ with parameter $n\in\mathbb{Z}_{\geq 0}$ by the equality
$r_{n,d}(x)=\left(-\frac{s^{2}}{n}\right)^{d}\cdot d!\cdot L_{d}^{(n)}\left(\frac{nx}{s^{2}}\right).$
Moreover, the unique monic polynomial of degree $d$ with $(n,d)$-rectangular cumulants being $K_{2\ell}^{n,d}=\delta_{\ell,1}$ is exactly $d!\cdot L_{d}^{(n)}(-x)$, as can be verified from (43). This is then the limiting object for the following rectangular finite free version of the Central Limit Theorem:
If $p(x)$ is any monic polynomial of degree $d$ with second cumulant $K_{2}^{n,d}$ equal to $1$, then the polynomials
$N^{-d}\cdot\big(p\boxplus_{d}^{n}\cdots\boxplus_{d}^{n}p\big)(Nx)\quad\text{(N-fold convolution)}$
converge coefficient-wise, as $N\to\infty$, to the polynomial $d!\cdot L_{d}^{(n)}(-x)$.
The symmetric finite free version of this result was stated and proved in [Mar-21, Thm. 6.7], and then revisited in [AP-18, Example 6.1], with the use of finite free cumulants. The rectangular version stated above can be proved by exactly the same argument as in the second reference, so we omit the details.
@@
In this section, [^7] Since $q$ has already appeared as the parameter for the theory of rectangular free probability, we use the bold $\mathbf{q}$ in this section to denote the “quantum” deformation parameter. we consider a $\mathbf{q}$-analogue of the Main Result I from the introduction; in order to state it, we need some classical notations from $\mathbf{q}$-analysis, e.g. from [GR-11]. The $\mathbf{q}$-numbers are $n]_{\mathbf{q}}=(1-\mathbf{q}^{n})/(1-\mathbf{q})$, $n\in\mathbb{Z}_{\geq 1}$, while $0]_{\mathbf{q}}=1$; the $\mathbf{q}$-factorial is $n!_{\mathbf{q}}=[1]_{\mathbf{q}}[2]_{\mathbf{q}}\cdots[n]_{\mathbf{q}}$, for $n\in\mathbb{Z}_{\geq 1}$, while $0!_{\mathbf{q}}=1$; the $q$-exponential function is $\exp_{\mathbf{q}}(x)=\sum_{n=0}^{\infty}{x^{n}/n!_{\mathbf{q}}}$.
@@
In the relations above, we used the $\mathbf{q}$-numbers $\ell]_{\mathbf{q}}=\frac{1-\mathbf{q}^{\ell}}{1-\mathbf{q}}$, $k]_{\mathbf{q}}=\frac{1-\mathbf{q}^{k}}{1-\mathbf{q}}$, but $\mathbf{q}^{-\frac{1}{2}}$-composition (namely, definitions (79)–(80) should be employed with the parameter $\mathbf{q}^{-\frac{1}{2}}$ instead of $\mathbf{q}$) for the expressions of the form $\exp_{\mathbf{q}^{-\frac{1}{2}}}[f(z)]$ on the left hand sides of (82)–(83).
@@
    where $g(z):=\sum_{\ell=1}^{\infty}{\kappa_{2\ell}z^{2\ell-1}}$; $T_{\mathbf{q}^{-\frac{1}{2}}}$ is the shift operator $T_{\mathbf{q}^{-\frac{1}{2}}}f(z):=f\big(\mathbf{q}^{-\frac{1}{2}}z\big)$; $z^{0}]f(z)$ is the constant term of the power series $f(z)\in\mathbb{R}[[z]]$; and finally, $\Delta_{\mathbf{q};t,u}$ is the linear operator on $\mathbb{R}[[z]]$ defined uniquely by its action on monomials:
@@
[…]
@@
$\begin{gathered}\partial_{\mathbf{q};t,u}z^{2n+1}:=\frac{\mathbf{q}^{-n-\frac{1}{2}}-u\mathbf{q}^{-\frac{1}{2}}}{\mathbf{q}^{-1}-1}z^{2n},\quad n\in\mathbb{Z}_{\geq 0},\\ \partial_{\mathbf{q};t,u}z^{2n}:=\frac{\mathbf{q}^{-n}-t}{\mathbf{q}^{-1}-1}z^{2n-1},\quad n\in\mathbb{Z}_{\geq 1},\qquad\partial_{\mathbf{q};t,u}(1):=0.\end{gathered}$ (87)